HUBBLE SPACE TELESCOPE DAILY REPORT #1542 PERIOD COVERED: 0000Z (UTC) 12/22/95 - 0000Z (UTC) 12/27/95 Daily Status Report as of 361/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of FOS/RD 6342 (The Spectrum of Background = Light) The Faint Object Spectrograph (Red Side) was used to make = spectrographic observations of back ground light. These observations were = taken in parallel during portions of the Hubble Deep Field (HDF) survey. = The data will be used to measure the extragalactic background light (EBL). = The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Two Sets of WF/PC-2 6337 (The Hubble Deep Field = Observations) The WF/PC-2 were used to take the Hubble Deep Field (HDF) images. = This project consists of hundreds of long WF/PC-2 exposures of the high = galactic latitude field (HDF-123649), taken in eight visits. A dark = calibration image was also taken. The observations were completed as = planned, and no problems were reported. 1.3 Completed Two Sets FOS/RD 6339 (Verification of the UV Spectrum of = the HST Background Light ) The Faint Object Spectrograph (Red Side) was used to make = ultraviolet observations of the HST background light. These FOS = observations were taken parallel to the Hubble Deep Field (HDF) survey. = Both the HDF and these FOS data will be used to test and improve the = existing engineering model of the HST background light. The observations = were completed as planned, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 6 Scheduled Reacquisitions : 33 Successful : 33 2.2 FHST Updates: Scheduled : 13 Successful : 13 2.3 Operations Notes: Temperatures aboard the HST are being closely monitored. = Excursions from normal ranges are being observed, principally in Bay #10, = due to the HST attitude being at the same position for the duration of the = multi-day Deep Field Survey and due to higher than nominal temperatures = that regularly occur during this season. Data analysis reveals that the Voltage Sensing Circuits (VSCs) = activated on days 349, 352, and 358 without explanation. These circuits = are designed to trigger when two or more VSCs sense that the diode bus = voltage has exceeded 32.7 volts. The VSCs then open relays that shunt an = additional isolation diode to decrease the diode bus voltage by = approximately 0.8 volts. Analysis of these VSC activations in under = investigation at this time. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH